# Discrete Optimal Transport and Voice Conversion

### v1

@@ Abstract
In this work, we address the voice conversion (VC) task using a vector-based interface. To align audio embeddings between speakers, we employ discrete optimal transport mapping. Our evaluation results demonstrate the high quality and effectiveness of this method. Additionally, we show that applying discrete optimal transport as a post-processing step in audio generation can lead to the incorrect classification of synthetic audio as real.
@@ 1 Introduction
Additionally, we explore the ability of the discrete OT to work on more distinct unpaired domains. We apply the OT-based mapping to convert generated audio samples into the domain of real speech using the ASVspoof 2019 dataset. As a result, in the majority of cases, the AASIST model [9] misclassified the converted audio as bona fide speech, indicating the effectiveness of the proposed approach (see Section 4.2).
@@ 2.1 Discrete Optimal Transport
The goal of optimal transport (OT) is to find the joint distribution $\pi\in\Pi(\mathbb{P},\mathbb{Q})$ known as Kantorovich plan or coupling, that minimizes the expected transport cost
@@
To compute the coupling $\gamma,$ we use entropic OT with Sinkhorn algorithm described in [10, Ch. 4].
@@ 2.2 Continuous Optimal Transport
Under mild conditions, the solution of the Kantorovich OT problem follows from the solution of the Monge problem [11, Th. 5.30].
When paired examples are available (e.g., from discrete OT), flow matching method can be used to find an approximate continuous transport map $T,$ as in [12, Th. 4.2].
2.3 Neural Optimal Transport
There is a rigorous mathematical proof that the OT problem 1 allows a minimax formulation with respect to two functions, which can be approximated with neural networks. One of those functions represents the transport map $T$ (see paper [2] for details, it also contains weak OT formulation, which allows to learn $\pi(y|x)$). Interestingly, the same research group later presented the OT without using neural networks, where the transport map is represented by a Gaussian mixture (see [13]).
@@ 3.4 Transportation Map
In the discrete OT approach, we compute the coupling matrix $\gamma$ for the marginal distributions 7 and cost function 8. For each $x_{i},$ we sort the target embeddings $y_{j}$ in decreasing order of $\gamma_{ij},$ denoting the sorted vectors as $y^{ot(i)}_{j}.$ The sorted coupling weights along each row (with fixes $i$) are denoted by $\gamma^{sort}_{ij}.$

### v2

@@ Abstract
In this work, we address the task of voice conversion (VC) using a vector-based interface. To align audio embeddings across speakers, we employ discrete optimal transport (OT) and approximate the transport map using the barycentric projection. Our evaluation demonstrates that this approach yields highquality and effective voice conversion. We also perform an ablation study on the number of embeddings used, extending previous work on simple averaging of kNN and OT results. Additionally, we show that applying discrete OT as a postprocessing step in audio generation can cause synthetic speech to be misclassified as real, revealing a novel and strong adversarial attack.
@@ 1 Introduction
Additionally, we explore the ability of the discrete OT to work on more distinct unpaired domains. We apply the OT-based mapping to convert generated audio samples into the domain of real speech using the ASVspoof 2019 dataset. As a result, in the majority of cases, the AASIST model [9] misclassified the converted audio as bona fide speech, indicating the effectiveness of the proposed approach for domain adoptation (see Section 4.2) and introducing a new adversarial attack.
@@ 2.1 Discrete Optimal Transport
The goal of optimal transport (OT) is to find the joint distribution $\pi\in\Pi(\mathbb{P},\mathbb{Q})$ known as Kantorovich plan, that minimizes the expected transport cost
@@
To compute the plan $\gamma,$ we use entropic OT with Sinkhorn algorithm described in [10, Ch. 4].
@@ 2.2 Continuous Optimal Transport
Under mild conditions, the solution of the Kantorovich OT problem follows from the solution of the Monge problem [11, Th. 5.30]. When paired examples are available (e.g., from discrete OT), flow matching method can be used to find an approximate continuous transport map $T,$ as in [12, Th. 4.2]. For neural network approximation of the transportation map see paper [2].
@@ 3.4 Transportation Map
In the discrete OT approach, we compute the matrix $\gamma$ for the marginal distributions 7 and cost function 8. For each $x_{i},$ we sort the target embeddings $y_{j}$ in decreasing order of $\gamma_{ij},$ denoting the sorted vectors as $y^{ot(i)}_{j}.$ The sorted weights along each row (with fixes $i$) are denoted by $\gamma^{sort}_{ij}.$